OpenStudy (anonymous):

2/3 of what number is 16?

OpenStudy (anonymous):

solve the equation x*(2/3)=16

OpenStudy (anonymous):

x=16*(3/2)=48/2=24

OpenStudy (anonymous):

24*2/3=8*2=16